The return on shares of Valley Transporter is predicted under the following various economic conditions: Recession-0.12 Normal +0.09 Boom +0.18 If each economy state has the same probability of occurring, what is the variance of the stock? Place your answer in decimal form using four decimal places.
Answer #1

Expected return=Respective return*Respective probability

=(1/3*-0.12)+(1/3*0.09)+(1/3*0.18)(probability of each=(1/3))

=0.05

probability Return probability*(Return-Expected Return)^2
1/3 -0.12 1/3*(-0.12-0.05)^2=0.0096333
1/3 0.09 1/3*(0.09-0.05)^2=0.0005333
1/3 0.18 1/3*(0.18-0.05)^2=0.0056333
Total=0.0158(Approx).

Standard deviation=[Total probability*(Return-Expected Return)^2/Total probability]^(1/2)

=0.1257(Approx).

Variance=Standard deviation^2

=0.0158(Approx).